OpenStudy (anonymous):

change 570 degrees to radian measure in terms of pie?

OpenStudy (anonymous):

divide by \(180\) and then stick a \(\pi\) next to it

OpenStudy (anonymous):

9.94?

OpenStudy (anonymous):

my guess is they want a fraction in reduced form

OpenStudy (anonymous):

but then its 497/50

OpenStudy (turingtest):

your why 497?decimal answer is (almost) correct. if you round you should get 9.95 radians. however if you want it in reduced fraction form you have\[570\cdot\frac\pi{180}=\frac{570\pi}{180}\]now reduce\[\frac{570}{180}\]

OpenStudy (turingtest):

why 497 ? your decimal answer is (almost) correct* ...typo

OpenStudy (anonymous):

but the terms of pie

OpenStudy (turingtest):

yeah, so don't mess with \(\pi\) since it can't be reduced just stick it in the denominator after reducing the fraction part

OpenStudy (anonymous):

570*3.14/180= 9.94

OpenStudy (turingtest):

well, first off, \[\pi\approx3.14159265259...\] and so your wounding is wrong I think, it rounds to 9.95, but that's still only an approcimate answer, and we almost always want *exact* answers in mathematics, so writing 9.95 is not the way to go. but the point is that satellite was telling you to write it in reduced fraction from do you remember how to reduce fractions like 2/4, and 8/32 ? reduce 570/180 and stick \(\pi\) next to it to get an exact answer

OpenStudy (turingtest):

rounding*

OpenStudy (turingtest):

i.e. just write the symbol \(\pi\) so the answer is exact

OpenStudy (anonymous):

9.95 π ?

OpenStudy (turingtest):

no, we want to *reduce* 570/180 do you remember how we can reduce like... 2/4=1/2 12/68=6/34=3/17 etc ? that's that you should do here. you

OpenStudy (turingtest):

you should not round off, as that makes your answer inexact

OpenStudy (turingtest):

if you cannot reduce 570/180, you need to review algebra

OpenStudy (anonymous):

i get u